# Interpretable Embeddings of Speech Enhance and Explain Brain Encoding Performance of Audio Models

### v1

@@ 2.4 Encoding Model Construction
To measure the performance of the encoding models, we used correlation (Pearson’s $r$) and the variance explained $R^{2}$, two standard metrics for evaluating encoding models (Vaidya, Jain, and Huth 2022; Anderson, Davis, and Lalor 2024; Goldstein et al. 2025; Tuckute et al. 2023; Li et al. 2023). Generally, we used Pearson’s $r$ to compare the encoding model performance of different representations, while the variance explained $R^{2}$ was used to quantify the unique and shared contributions of different features for neural prediction. Negative $R^{2}$ values were clipped to 0, since negative explained variance lacks interpretability.
Variance Partitioning
@@
To quantify the shared and unique variance contributions of feature sets, we used the variance partitioning method established by de Heer et al. (Heer et al. 2017). For details, refer to Section S4 or Heer et al. 2017.
Computational Resources
@@ 3 Results
3.1 Interpretable models outperform and enhance SSMs in predicting brain responses
Figure 2: Interpretable model outperforms and enhances SSM representations in predicting brain responses. (a) Prediction performance of the model based on our interpretable features for all electrodes, plotted on a template brain. (b) Average Pearson’s $r$ across all electrodes for different models. For each SSM, the context size with the highest average Pearson’s $r$ was chosen. Joint Model refers to a combination of interpretable features and Whisper 32nd layer representations with 30 seconds of context. The interpretable model outperformed all SSMs, and the joint model of interpretable features and SSM embeddings achieved the highest average Pearson’s $r$. (c) Pearson’s $r$ for different models with the best layer and context size, shown for different anatomical regions (Paired t-test, FDR correction, two-sided). The interpretable model performs well across anatomical regions and significantly outperforms SSMs in lower-level regions, such as Heschl’s gyrus and the insula.
Fig. 2a shows the spatial distribution of electrode-wise correlation values for our interpretable model, mapped onto a template brain. Fig. S1 quantifies the contribution of unique variance explained $R^{2}$ by each feature set. Notably, the language processing hierarchy can be clearly observed in the anatomical distributions.
Fig. 2b shows the average Pearson’s $r$ values across all electrodes for each SSM over layers, compared with our interpretable model. For each model, the context size with the highest peak average correlation value was chosen and noted. Our interpretable model achieved an average correlation value of 0.079, outperforming the highest average correlation values of 0.073, 0.073, and 0.069 observed for WavLM, Whisper, and HuBERT, respectively ($p<0.001$, paired t-test, Fig. S2). The joint models of interpretable features with SSM representations achieved average correlation values of approximately 0.082 (0.0828 for WavLM, 0.0827 for Whisper, and 0.0821 for HuBERT), significantly outperforming models using only interpretable features ($p<0.001$, paired t-test). In agreement with the previous findings (Vaidya, Jain, and Huth 2022; Tuckute et al. 2023; Anderson, Davis, and Lalor 2024), all three models significantly outperformed the model based solely on mel-spectrogram, which achieved an average correlation of 0.057 ($p<0.001$, paired t-test). An anatomical analysis of encoding performance (Fig. 2c) shows that the interpretable model consistently outperforms all SSMs in Heschl’s gyrus and the insula—two regions known to encode low-level acoustic features (Berger et al. 2025; Kurteff et al. 2024; Khalighinejad et al. 2021; Leaver and Rauschecker 2010). Even when the comparison is applied to early layers of SSMs, this difference remains statistically significant, except for HuBERT, although the gap in the average correlation slightly narrows (Fig. S3). These results match with findings from Li et al. that concluded that HG representation is not modeled well by speech DNNs (Li et al. 2023).
Previous work showed that later layers of self-supervised models (SSMs) discard low-level acoustic details (Pasad, Chou, and Livescu 2022; Pasad et al. 2024). If the discarded details are important for predicting neural responses, combining early and late layer representations should improve predictive performance. To test this hypothesis, we evaluated joint models of representations from early and late SSM layers (Fig. 2b). Due to the high dimensionality of the combined features, we varied $\alpha$ across 40 logarithmically spaced values from $10^{0}$ to $10^{6}$. Combining Whisper layers 1 and 32 improved average Pearson’s r by 0.003 over layer 32 alone; WavLM layers 1 and 20 improved by 0.00024; and HuBERT layers 1 and 32 by 0.0025. While these gains may be diminished by overfitting due to increased dimensionality, these gains remain substantially smaller than the predictive performance of the mel-spectrogram (r = 0.056), which can be viewed as an approximate upper bound on the expected improvement if later SSM layers had entirely discarded low-level acoustic information. This suggests that later SSM layers may retain more low-level acoustic information than previously assumed, at least in forms relevant for neural processing.
3.2 SS Models compress and retain brain-relevant low-level acoustic information
Figure 3: SSMs compress and retain brain-relevant low-level acoustic information. (a) The ratio of neural variance shared by the mel-spectrogram and the SSM representations, relative to the total variance explained by the mel-spectrogram. The ratio remains consistently high across layers, suggesting that acoustic information relevant for neural processing is preserved. (b) Average Variance $R^{2}$ of the SSM representations across their feature dimensions, explained by the mel-spectrogram. This metric quantifies how much variance in SSM embeddings is allocated to encoding mel-spectrogram information, and shows a substantial decrease in later layers. (c) Variance $R^{2}$ of the mel-spectrogram [$t-0.5$, $t$]s explained by the Whisper representations over different layers. (d) Average variance explained $R^{2}$ over the 0.2 seconds [$t-0.2$, $t$]s for different frequency bins, plotted over layers. The gain/loss in percentage with respect to the 1st layer is displayed on the middle plot. On the right plot, the gain/loss percentage for the 32nd layer is shown, demonstrating the preferential preservation of speech critical frequency bands.
Inspired by the findings in the previous section, we conducted a mechanistic interpretability analysis on how low-level acoustic information is represented across the layers of the SSMs.
Fig. 3a shows the proportion of neural variance in Heschl’s gyrus and Insula shared by the mel-spectrogram and the SSM representations, relative to the total variance explained by the mel-spectrogram. This ratio reflects the extent to which each layer of an SSM encodes brain-relevant low-level acoustic information. As shown in the figure, this ratio remains consistently high across layers: for Whisper, it peaks at 90.4% and decreases to 77.5% in later layers; for WavLM, from 89.5% to 80.1%; and for HuBERT, from 96.2% to 71.9%.
To better understand how the mel-spectrogram is represented internally within each model, Fig. 3b shows the average variance explained by the mel-spectrogram when directly predicting SSM representations across layers. This metric quantifies the degree to which the internal representations of the SSMs are allocated to encode the feature. For Whisper, the average explained variance $R^{2}$ peaks at layer 1 with a value of 0.204 and drops to 0.093 in the last layer, corresponding to relative reductions of 54.4%. For WavLM, it decreases from 0.192 to 0.120 (37.5% reduction); and for HuBERT, from 0.234 to 0.055 (76.5% reduction). Notably, these reductions are substantially larger than the relatively modest declines observed in neural variance shared between the mel-spectrogram and SSM representations in Heschl’s gyrus and Insula, suggesting that while low-level acoustic information becomes increasingly compressed within the models, it remains selectively preserved in brain-relevant representations.
Next, to investigate how these models retain brain-relevant low-level acoustic information over layers, we directly predicted 0.5-second windows of mel-spectrogram from the SSM representations: that is, using the SSM representations at time t s, we predicted the entire mel-spectrogram in the [$t-0.5$, $t$]s window. Fig. 3(c) shows the variance $R^{2}$ of the mel-spectrogram explained by Whisper representations over different layers. It can be observed that in the earlier layers (0th and 8th), the Whisper representations predict all frequency bins equally well, but their predictive performance in the lowest frequency bin and higher frequency bins drops markedly in later layers. Fig. 3(d) quantifies this trend by averaging the variance explained across the [$t-0.2$, $t$]s window for each frequency bin. Frequencies in the 100-1000 Hz range exhibit a significantly smaller decrease in explained variance compared to frequencies outside this range (average: $9.9$% vs. $17.9$%). Fig. S5 shows a similar trend for WavLM and HuBERT as well. This frequency band aligns with regions of greater cortical representation ($<$ 1000 Hz) during speech perception, as previously reported in Heschl’s gyrus and surrounding areas (Moerel, Martino, and Formisano 2012; Khalighinejad et al. 2021). Together, these results suggest that while SSMs compress low-level acoustic features internally, they preferentially preserve the components most relevant to brain responses during speech perception.
3.3 SS Models Encode Brain-Relevant Semantic Information
@@
Figure 4: SSMs encode brain-relevant semantic information. (a) Diagram of the three-way variance partitioning analysis used to assess whether SSM representations encode semantic information. (b) Unique variance explained $R^{2}$ by different feature sets in predicting GPT-2 representations, averaged across GPT-2’s feature dimensions ($p<0.001$, Wilcoxon signed-rank test). The significant unique contribution of our question-answering semantic features confirms that they encode distinct semantic information. (c) Unique neural variance explained by semantic QA features relative to non-semantic interpretable features. (d) Box plot of the values in (c), across electrodes in the inferior frontal gyrus triangularis (IFG-Tri). Median values were significantly greater than zero ($p<0.001$, Wilcoxon signed-rank test). (e) $\Delta R^{2}$, as illustrated in (a), averaged across electrodes in IFG-Tri. Left: Varying context lengths, using the largest variants of each SSM. Right: Varying model sizes, with context length fixed to 5 seconds for WavLM and HuBERT, and 30 seconds for Whisper. For each model size and context length, the layer with the highest $\Delta R^{2}$ was chosen. It shows SSMs encode brain-relevant semantic information that cannot be reduced to lower-level features, improving with context length and model size.
@@
Fig. 4a outlines our approach for evaluating whether SSM representations encode semantic information. First, we calculate the unique neural variance $R^{2}$ explained by semantic QA features with respect to non-semantic interpretable feature sets (mel-spectrogram, GBFB, speech presence, phonetic, and syntactic features). We then assess how much this unique neural variance is reduced when SSM representations are included in the model alongside the non-semantic features. This reduction reflects the shared neural variance $R^{2}$ between semantic QA features and SSM representations that is not explained by non-semantic interpretable feature sets.
QA features encode genuinely semantic information relevant for brain responses
@@
Before we assess SSMs, to evaluate whether our QA features actually encode semantic information, we computed the unique variance $R^{2}$ by each interpretable feature set in predicting GPT-2 representations (Fig. 4b). This step was included to ensure the quality of QA features, which depend on the accuracy of the LLM’s answers. Syntactic, semantic, and phonetic features account for the largest proportion of unique variance ($p<0.001$, Wilcoxon signed-rank test) across GPT-2’s feature dimensions. Speech probability also explains a statistically significant amount of unique variance, though its contribution is markedly smaller in comparison. In contrast, GBFB and mel-spectrogram features do not explain any significant unique variance, as expected—GPT-2 is not trained on acoustic input and thus should not encode such information. The significant unique variance explained by our semantic features supports the claim that they indeed capture meaningful semantic content.
@@
Next, to test whether semantic information captured by QA features is relevant to neural processing, we computed the variance of the brain responses uniquely explained by semantic features relative to non-semantic interpretable features (Fig. 4c). Semantic features had the largest unique variance explained in the inferior frontal gyrus pars triangularis (IFG-Tri) - a part of Broca’s area that was previously found to be important for processing the semantic meaning of words (Newman et al. 2003; Vigneau et al. 2006; Friederici, Opitz, and von Cramon 2000). As shown in Fig. 4(d), the median of the unique neural variance explained by semantics in IFG-Tri was significantly above 0 ($p<0.001$, Wilcoxon signed-rank test). These results confirm that our semantic QA features are not only semantically meaningful but also relevant to brain regions involved in semantic comprehension.
SS models encode semantic information that is relevant for brain responses
@@
The left plot of Fig. 4e shows $\Delta R^{2}$, the shared neural variance between semantic QA features and SSM representations that is not accounted for by non-semantic interpretable features, across different context sizes. Averaged across all electrodes in IFG-Tri, $\Delta R^{2}$ was significantly above 0 ($p<0.001$, Wilcoxon signed-rank test). For WavLM and HuBERT, it increases up to a 5-second context window, consistent with the 4.75-second intrinsic context of our semantic QA features. For Whisper, the increase extended up to 30 seconds, potentially because Whisper was trained strictly with 30 seconds of audio. The right plot in Fig. 4e shows the same metric, but for varying model sizes with fixed context size. Here, the average $\Delta R^{2}$ increases with model size, suggesting that larger models encode more brain-relevant semantic information.

### v2

@@ 2.4.2 Performance Metrics
To measure the performance of the encoding models, we used correlation (Pearson’s $r$) and the variance explained $R^{2}$, two standard metrics for evaluating encoding models (Vaidya et al., 2022; Anderson et al., 2024; Goldstein et al., 2025; Tuckute et al., 2023; Li et al., 2023). Generally, we used Pearson’s $r$ to compare the encoding model performance, while the variance explained $R^{2}$ was used to quantify the unique and shared contributions of different features for neural prediction. Negative $R^{2}$ values were clipped to 0, since negative explained variance lacks interpretability, following Tuckute et al. (2023); Hadidi et al. (2025); Anderson et al. (2024).
2.4.3 Variance Partitioning
@@
To quantify the shared and unique variance contributions of feature sets, we used the variance partitioning method established by Heer et al. (2017). Between two feature sets $\mathcal{A}$ and $\mathcal{B}$,
$R^{2}_{\text{unique}(\mathcal{A})}=R^{2}_{\mathcal{A},\mathcal{B}}-R^{2}_{\mathcal{B}}\qquad R^{2}_{\text{shared}(\mathcal{A},\mathcal{B})}=R^{2}_{\mathcal{A}}+R^{2}_{\mathcal{B}}-R^{2}_{\mathcal{A},\mathcal{B}}$
where $R^{2}_{\mathcal{A}}$, $R^{2}_{\mathcal{B}}$, and $R^{2}_{{\mathcal{A},\mathcal{B}}}$ denote the variance explained by models using only $\mathcal{A}$, only $\mathcal{B}$, and both $\mathcal{A}$ and $\mathcal{B}$ together, respectively. All models were fitted using banded ridge regression.
2.4.4 Computational Resources
@@ 3 Results
Figure 2: Decomposing SFM Brain Encoding Performance with Interpretable Features. (a) Diagram of the variance partitioning analysis used to answer three key questions about the encoding model performance of SFMs. (b) The ratio of neural variance shared by the SFM representations and the combined set of all interpretable features, relative to the total variance explained by the SFM representations. In early layers, the interpretable model explains nearly all of the neural variance, and even in later layers, it continues to account for a substantial majority. (c) The ratio of neural variance shared between different interpretable feature sets and the SFM representations, relative to the total variance explained by each interpretable feature set. This ratio peaks in early layers for the mel-spectrogram, but in progressively later layers for higher-level features (phonetic, syntactic, and semantic). (d) Unique neural variance of SFMs with respect to the combined set of all interpretable features, averaged across all electrodes. For all three models, the increase is monotonic over layers. (e) Unique neural variance of SFMs as a function of model size. For each model size, the layer with the highest unique variance was selected. Results show that the uniquely explained neural component increases with model size. (f) Unique neural variance explained by the embeddings from Whisper layer 32 with respect to the combined set of all interpretable features, plotted on a template brain. The largest unique contributions appear in the superior temporal gyrus (STG) and inferior frontal gyrus (IFG), regions that are typically associated with the processing of higher-order information.
3.1 Decomposing SFM Brain Encoding Performance with Interpretable Features
Fig.˜2 decomposes SFM encoding model performance into variance explained by interpretable features versus unexplained by them, and examines this decomposition from three perspectives (Fig.˜2a): (i) To what extent can the neural variance explained by SFMs be attributed to interpretable features? (ii) How does the shared neural variance between SFMs and individual feature sets change across model layers? (iii) If there is any neural variance that cannot be attributed to known interpretable features, how does it evolve across model layers? Does it scale with the number of model parameters?
Most of the neural variance captured by SFMs is attributable to interpretable features. Fig.˜2b shows the ratio of neural variance shared between the SFM representations and a combined set of all interpretable features, relative to the total variance explained by the SFM representations. This metric directly quantifies what proportion of the variance captured by SFMs can be attributed to known, interpretable features. This ratio is particularly high in early layers, exceeding 95% in the earliest layers of these models. Even among the latest layers, these proportions remained substantially high at above 75% respectively. These results suggest that in early layers, nearly all the neural variance captured by SFMs can be explained by interpretable features. Although the ratio decreases in later layers, it stays substantial, implying that interpretable confounds continue to explain much of the neural variance.
SFMs display a layerwise tradeoff between encoding of brain-relevant low-level and high-level features. Fig.˜2c displays the ratio of neural variance shared between each interpretable feature set and the SFM representations, relative to the total variance explained by that feature set alone. This ratio indicates how well the speech model captures the neurally relevant information contained in that feature space. Here, we report results for the mel-spectrogram, phonetic, syntactic, and semantic features; results for additional feature sets are provided in Fig.˜S4.
The ratios exhibited distinct trajectories for low-level versus high-level features. For the mel-spectrogram, the ratio peaked in early layers—at layer 0 for HuBERT (0.90) and WavLM (0.80), and at layer 5 for Whisper (0.79). In later layers, the ratios decreased, reaching the lowest values of 0.61, 0.73, and 0.71, respectively. While the decrease was modest for WavLM and Whisper, further analysis in Section˜3.3 suggests that this reflects a trade-off: although their embeddings lose precision, they concurrently gain broader temporal windows of acoustic information. A similar decreasing trend was observed for other low-level features, such as GBFB features and speech probability (see Fig.˜S4). In contrast, alignment with higher-order features emerged progressively across layers. For phonetic features, all three models showed a sharp increase after the initial layers, reaching near-maximal ratios (0.95 for HuBERT, 0.96 for WavLM, and 0.95 for Whisper) in middle-to-late layers, where they remained consistently high thereafter. Syntactic and semantic features exhibited a more gradual rise, with the ratios increasing steadily across layer depth and peaking in later layers than phonetic features. Together, these results indicate a trade-off in the neurally relevant components of SFM representations: gains in high-level representations emerge at the expense of lower-level features, rather than through a simple monotonic accumulation of information.
Unexplained neural variance of SFMs grows with layer depth and scale. Fig.˜2d shows the unique neural variance explained by the SFMs relative to the combined set of all interpretable features, averaged across electrodes. This unique variance increased monotonically across layers in all three models. For HuBERT, it peaked in the late-middle layers and remained high thereafter, whereas for WavLM and Whisper, the peak occurred in the final layers. This shows that the later layers of SFMs are the most useful for capturing components of neural variance that cannot be explained by interpretable counterparts. Fig.˜2e shows the average unique neural variance of SFMs as a function of model size. Unique variance increased systematically with the number of model parameters, indicating that the unexplained neural signal scales with model size. This result aligns with Antonello et al. (2023), who found that the total neural variance explained by SFMs increases logarithmically with the number of parameters.
Fig.˜2f shows the unique neural variance explained by the final layer of Whisper, relative to the combined set of all interpretable features, projected onto a template brain. The largest unique contributions appear in the superior temporal gyrus (STG) and inferior frontal gyrus (IFG), regions that are typically associated with processing of higher-order information (Newman et al., 2003; Vigneau et al., 2006; Friederici et al., 2000; Keshishian et al., 2023). Similar spatial patterns are observed for HuBERT and WavLM, as presented in Fig.˜S5. Combined with earlier findings that most neural variance in the early layers can be explained by interpretable features, these results suggest that the remaining unique variance in later layers likely reflects high-level, non-linear representations that are not captured by the current interpretable feature sets.
Figure 3: Interpretable features efficiently enhance the encoding model performance of SFMs. Test correlation value for each SFM and its joint model with the combined set of all interpretable features, averaged across electrodes. Left: WavLM, Middle: Whisper, Right: HuBERT. Combining interpretable features with SFMs offers two key advantages: (1) it compensates for the fact that no single SFM layer captures all information relevant for neural processing, and (2) it leverages the increasing amount of unexplained variance in deeper SFM layers, which also scales with model size. Results show that incorporating interpretable features consistently improves encoding performance across all layers (p < 0.001, Wilcoxon signed-rank test).
3.2 Building More Interpretable and Accurate Brain Encoding Models by Augmenting SFM representations with Interpretable Features
Our findings in the previous section reveal two key insights. First, there is a trade-off between low-level and high-level feature encoding in SFM representations across layers. This suggests that no single layer contains all the neural-relevant features necessary for modeling speech perception. Second, deeper layers account for additional neural variance that is not captured by our interpretable features, and this unique contribution increases with model size. These observations motivate a principled approach for building more interpretable and accurate encoding models: combining interpretable feature sets with the later layers of large SFMs. This strategy mitigates the limitations of relying on individual layers while harnessing the representational advantages conferred by scaling, resulting in a more comprehensive and effective neural encoding model for speech.
Fig.˜3 shows the average Pearson’s $r$ for models based on SFM representations alone and for the joint model combining SFM representations with interpretable features. For all models, incorporating interpretable features consistently improves encoding performance across all layers (p < 0.001, Wilcoxon signed-rank test). Importantly, the joint models also simplify interpretation. For WavLM and HuBERT, correlations based solely on SFM representations peak in mid-to-late layers but then decline toward the final layers. When interpretable features are incorporated, this drop-off is substantially reduced as the information lost in the final layers is encoded by the interpretable feature sets. In the joint model, encoding performance always reflects contributions from interpretable features spanning low-level acoustic to high-level lexical domains, while the unique neural variance explained by the SFM representations grows with depth. This results in a more interpretable and accurate model of the brain.
Figure 4: Whisper loses fine-acoustic details and gains longer acoustic temporal windows in later layers. (a) We quantify the shared neural variance between Whisper representations and mel-spectrogram features with different window lengths. (b) Average neural variance $R^{2}$ shared between Whisper representations and mel-spectrogram segments of different temporal windows. Across all anatomical regions, shared variance with the short window [t0.05t-0.05, $t$]s consistently decreases across layers, whereas shared variance with the longer window [t0.5t-0.5, $t$]s exhibits anatomically dependent patterns. (c) Variance $R^{2}$ of the mel-spectrogram [t0.5t-0.5, $t$]s explained by the Whisper representations over different layers. In early layers, a high proportion of variance is explained for short windows, while in later layers, the explained variance is lower but reflects information from broader temporal windows.
3.3 Architecture-Dependent Trade-off Between Precision and Temporal Windows of Acoustic Information Encoded by SFMs
Our earlier finding in Section˜3.1—that the neural variance shared between the mel-spectrogram and Whisper or WavLM decreases only modestly across layers—is puzzling at first sight. This stability contrasts with evidence from Pasad et al. (2022; 2023) indicating that speech foundation models progressively lose fine acoustic details as information propagates through their transformer layers. To reconcile this apparent discrepancy, we tested a new hypothesis: the last token embeddings of Whisper and WavLM encode short, precise windows of mel-spectrogram in early layers, but shift toward encoding longer windows with less precision in later layers, maintaining the overall shared variance.
To test this hypothesis, we quantify the shared neural variance between Whisper representations and mel-spectrogram features with different window lengths (Fig.˜4a). Here, we present results for Whisper; results for WavLM and HuBERT are provided in Fig.˜S6. Fig.˜4b shows the average shared neural variance $R^{2}$ between Whisper representations and mel-spectrograms computed using four temporal windows (0.05, 0.1, 0.3, and 0.5 seconds), across the whole brain as well as in Heschl’s gyrus (HG), the Insula, and the superior temporal gyrus (STG). Importantly, the gap between each successive lines shows changes in shared variance as the window length increases, quantifying how much Whisper representations capture the neural variance uniquely attributable to mel-spectrogram information within specific temporal windows. As shown in the plot, the shared variance decreases monotonically with the shortest window ([0, 0.05]s) and increases with longer windows ([0.1, 0.5]s), suggesting that Whisper embeddings progressively lose fine, short-timescale acoustic detail while gaining longer-timescale acoustic information.
Notably, different anatomical regions display distinct trends. In HG and the Insula, the shared variance decreases across layers even when using the 0.5-second mel-spectrogram, whereas in STG, it increases over layers until it plateaus at around layer 16. For STG, the gain in shared variance with longer windows ([0.1, 0.5]s) outweighs the loss observed with shorter ones ([0, 0.05]s), while the opposite is true for HG and the Insula. This likely reflects differences in their temporal integration windows, which have been shown to lengthen from primary to non-primary auditory areas (Lerner et al., 2011; Sabat et al., 2025; Norman-Haignere et al., 2022). HG, with its short integration window, is sensitive to fine acoustic detail, while STG, with its longer integration window, prioritizes broader temporal patterns.
To further test our hypothesis, we directly predicted 0.5-second segments of the mel-spectrogram from the SFM representations. Specifically, we used the representations at time $t$ to predict the corresponding mel-spectrogram within the [t0.5t-0.5, $t$]s window. Fig.˜4c shows the variance $R^{2}$ of mel-spectrogram explained by Whisper representations across different layers. In the 0th layer, Whisper representations accurately predict all frequency bins for the shorter segments of the mel-spectrogram. However, as information progresses through the network, the explained variance for these short segments decreases. At the same time, the predictions become temporally broader, indicating that deeper layers shift toward encoding longer-range temporal structure at the expense of fine-grained acoustic detail. This pattern provides further support for the hypothesis that the last token embeddings of Whisper encode short, precise windows of mel-spectrogram in early layers, but shift toward encoding longer windows with less precision in later layers.
Similar trends were observed for WavLM (see Fig.˜S6a-b). In contrast, HuBERT (Fig.˜S6c-d) showed a qualitatively different pattern: already at layer 0, its embeddings captured the mel-spectrogram information in longer windows (0.1-0.5 s), with little further increase between layers. These differences likely arise from how positional information is incorporated prior to the transformer layers. Both HuBERT and WavLM use a convolutional layer to add relative positional embeddings, extending the temporal receptive window of a token beyond that of the CNN encoder. WavLM further applies a gating mechanism on top of this convolution, which may explain its divergence from HuBERT. Whisper, in contrast, applies fixed sinusoidal positional embeddings directly to the log-mel spectrogram, constraining receptive windows of a token to those of the spectrogram itself. As a result, when analyzing last-token embeddings as done in Vaidya et al. (2022); Antonello et al. (2023); Oota et al. (2024), Whisper and WavLM exhibit a clear layer-wise trade-off—progressively shifting from fine-grained, short-timescale acoustic detail to broader temporal integration—whereas HuBERT’s acoustic representations simply degrade in later layers. Together, these findings clarify why Whisper and WavLM maintain stable shared variance across layers: not by preserving detail, but by broadening their temporal receptive windows. More broadly, they highlight how architectural choices shape the dynamics of acoustic information across model layers.
@@ 3.1 Decomposing SFM Brain Encoding Performance with Interpretable Features
Figure 5: SFMs encode brain-relevant semantic information. (a) Diagram of the three-way variance partitioning analysis used to assess whether SFM representations encode semantic information. (b) Unique variance explained $R^{2}$ by different feature sets in predicting GPT-2 representations, averaged across GPT-2’s feature dimensions ($p<0.001$, Wilcoxon signed-rank test). The significant unique contribution of our question-answering semantic features confirms that they encode distinct semantic information. (c) Unique neural variance explained by semantic QA features relative to non-semantic interpretable features. (d) Box plot of the values in (c), across electrodes in the inferior frontal gyrus triangularis (IFG-Tri). Median values were significantly greater than zero ($p<0.001$, Wilcoxon signed-rank test). (e) $\Delta R^{2}$, averaged across electrodes in IFG-Tri. Left: Varying context lengths, using the largest variant of each SFM. Right: Varying model sizes, with context length fixed to 5 seconds for WavLM and HuBERT, and 30 seconds for Whisper. It shows SFMs encode brain-relevant semantic information that cannot be reduced to lower-level features, improving with context length and model size.
3.4 SFMs Encode Brain-Relevant Semantic Information
@@
Fig.˜5a outlines our approach for evaluating the degree to which SFM representations encode brain-relevant semantic information. First, we calculate the unique neural variance $R^{2}$ explained by semantic QA features with respect to non-semantic interpretable feature sets (mel-spectrogram, GBFB, speech presence, phonetic, and syntactic features). We then assess how much this unique neural variance is reduced when SFM representations are included in the model alongside the non-semantic features. This reduction reflects the shared neural variance $\Delta R^{2}$ between semantic QA features and SFM representations that is not explained by non-semantic interpretable feature sets.
@@
Before we assess SFMs, to evaluate whether our QA features actually encode semantic information, we computed the unique variance $R^{2}$ by each interpretable feature set in predicting GPT-2 representations (Fig.˜5b). This step was included to ensure the quality of QA features, which depend on the accuracy of the LLM’s answers. Syntactic, semantic, and phonetic features account for the largest proportion of unique variance ($p<0.001$, Wilcoxon signed-rank test) across GPT-2’s feature dimensions. Speech probability also explains a statistically significant amount of unique variance, though its contribution is markedly smaller in comparison. In contrast, GBFB and mel-spectrogram features do not explain any significant unique variance, as expected—GPT-2 is not trained on acoustic input and thus should not encode such information. The significant unique variance explained by our semantic features supports the claim that they indeed capture meaningful semantic content.
@@
Next, to test whether semantic information captured by QA features is relevant to neural processing, we computed the variance of the brain responses uniquely explained by semantic features relative to non-semantic interpretable features (Fig.˜5c). Semantic features had the largest unique variance explained in the inferior frontal gyrus pars triangularis (IFG-Tri) - a part of Broca’s area that was previously found to be important for processing the semantic meaning of words (Newman et al., 2003; Vigneau et al., 2006; Friederici et al., 2000). As shown in Fig.˜5(d), the median of the unique neural variance explained by semantics in IFG-Tri was significantly above 0 ($p<0.001$, Wilcoxon signed-rank test). These results confirm that our semantic QA features are not only semantically meaningful but also relevant to brain regions involved in semantic comprehension.
@@
The left plot of Fig.˜5e shows $\Delta R^{2}$, the shared neural variance between semantic QA features and SFM representations that is not accounted for by non-semantic interpretable features, across different context sizes. Averaged across all electrodes in IFG-Tri, $\Delta R^{2}$ was significantly above 0 ($p<0.001$, Wilcoxon signed-rank test). For WavLM and HuBERT, it increases up to a 5-second context window, consistent with the 4.75-second intrinsic context of our semantic QA features. For Whisper, the increase extended up to 30 seconds, potentially because Whisper was trained strictly with 30 seconds of audio. The right plot in Fig.˜5e shows the same metric, but for varying model sizes with fixed context size. Here, the average $\Delta R^{2}$ increases with model size, suggesting that larger models encode more brain-relevant semantic information.